OpenStudy (matimaticas):

What is the axis of symmetry for f(x) = 2x^2 − 8x + 8?

OpenStudy (raden):

are you familiar with the formula : x = -b/(2a) ?

OpenStudy (matimaticas):

yeah

OpenStudy (s3a):

1) The axis of symmetry of f(x) = 2x^2 - 8x + 8 is the same as the axis of symmetry of 2x^2 - 8x (because it's just shifted vertically). 2) Find the zeroes of f(x) = 2x^2 - 8x. 0 = 2x^2 - 8x 0 = x(2x - 8) x = 0, 2x - 8 = 0 x = 0, x = 4 Horizontal component of vertex = (4 - 0)/2 = 2. Therefore, the axis of symmetry is the line x = 2.

OpenStudy (s3a):

(Because of the symmetry of the parabola, the horizontal component of the vertex is in the middle of the two zeroes.)

OpenStudy (matimaticas):

thanks

OpenStudy (s3a):

You're welcome. :)